# Surgical Gaussian Surfels: Highly Accurate Real-time Surgical Scene Rendering

### v1

@@ 3.2 PIMI Gaussian Surfel Primitive Initialization
To mitigate these issues, we propose an initialization method, termed Projection-Based Iterative Mask Integration (PIMI), which improves the quality and completeness of the initial point cloud by leveraging confidence-weighted depth aggregation and aggregated color mapping. The depth map $D_{t}$ for each frame is pre-processed to create a confidence-weighted depth map $D^{*}$ using a thresholded confidence mask, which ensures the exclusion of noisy depth data. Simultaneously, we aggregate the color maps $C_{t}$ to compute the refined color map $C^{*}$. This aggregation scheme at timestep $t$ is defined as:
@@
where $W_{t}$ is the confidence mask derived by identifying reliable depth values within the 2nd to 99th percentile range of the depth distribution. Using $D^{*}$ and $C^{*}$, we generate the refined point cloud $P^{*}$. Each 2D pixel coordinate $(x,y)$ is projected into 3D space based on the intrinsic camera parameters $K$ as:
@@ 3.3 Optimization
To capture fine details of surgical tissue and render areas occluded by surgical tools, we invert the binary mask in occluded areas and employ a weighted total loss consisting of photometric loss [14], total variation (TV) loss [11], deformation regularization, VGG16 [15] perceptual loss, and normal consistency:
$L=\lambda_{\text{1}}L_{\text{photo}}+\lambda_{\text{2}}L_{\text{tv}}+\lambda_{\text{3}}L_{\text{pos}}+\lambda_{\text{4}}L_{\text{cov}}+\lambda_{\text{5}}L_{\text{per}}+\lambda_{\text{6}}L_{\text{n}}.$ (6)
Deformation regularization. In single-view surgical video, surfel motion can become unstable without constraints. To ensure coherent yet flexible deformation, we use a deformation regularization that maintains relative distances between neighboring surfels before and after deformation. Given a surfel $\psi_{k}$, we enforce consistency across its $K$ nearest neighbors (where $K=5$[17] in both position and scaling:
$L_{\text{pos}}=\sum_{k}\sum_{j\in\mathcal{N}(k)}\left\|d(\mathbf{x}^{(k)},\mathbf{x}^{(j)})-d(\hat{\mathbf{x}}^{(k)},\hat{\mathbf{x}}^{(j)})\right\|_{1},$ (7)
$L_{\text{cov}}=\sum_{k}\sum_{j\in\mathcal{N}(k)}\left\|d(\mathbf{s}^{(k)},\mathbf{s}^{(j)})-d(\hat{\mathbf{s}}^{(k)},\hat{\mathbf{s}}^{(j)})\right\|_{1}.$ (8)
where $d(\cdot,\cdot)$ computes the Euclidean distance, and $\mathbf{x},\hat{\mathbf{x}}$ and $\mathbf{s},\hat{\mathbf{s}}$ denote initial and deformed surfel positions and scaling factors, respectively.
Normal consistency. The normal consistency term enforces alignment between the rendered normals and the implicit normals derived from the geometric properties of the scene. This regularization ensures that all Gaussian surfels are locally aligned with the tissue surfaces. The loss is defined as:
$L_{\text{n}}=\alpha\cdot(1-\cos(\mathbf{n}_{\text{rendered}},\mathbf{n}_{\text{expected}}))+\beta\cdot(1-\cos(\mathbf{n}_{\text{rendered}},\mathbf{n}_{\text{median}})),$ (9)
where $\mathbf{n}_{\text{rendered}}$, $\mathbf{n}_{\text{expected}}$, and $\mathbf{n}_{\text{median}}$ are the rendered normals, the expected normals from surface geometry (i.e., gradient of the depth map), and the normals of surfels at the median intersection point along the viewing ray, respectively. The weights $\alpha$ and $\beta$ balance the contributions of the two alignment targets.
Table 1: Quantitative comparisons of our method SGS with EndoNeRF [16], EndoSurf [20], LerPlane [18], EndoGaussian [10], and SurgicalGaussian (SG) [17].
Dataset Methods "pulling" "cutting" Rendering
@@
Ours 0.027 39.059 0.971 0.027 37.609 0.965 66 2 GB
"intestine" "liver" FPS↑ GPU↓
@@
Ours 0.048 34.669 0.936 0.047 34.390 0.934 81 2 GB
@@
w/o Perceptual loss 0.028 38.911 0.970 0.032 37.307 0.962
w/o Photometric loss 0.028 38.852 0.970 0.043 35.566 0.957
w GIDM initialization 0.027 38.907 0.970 0.031 37.380 0.963
Full model (PIMI init) 0.027 39.059 0.971 0.027 37.609 0.965
4 Experiments
Datasets and evaluation. We evaluate our method on two publicly available datasets, EndoNeRF [16] and StereoMIS [5]. The EndoNeRF dataset captures two cases of an in-vivo DaVinci robot prostatectomy from a single viewpoint. We evaluate on four video sequences of the StereoMIS dataset, which shows porcine subject breathing and tissue deformation during a liver and intestinal procedure. We estimate scene depth on the StereoMIS dataset using a pre-trained Omnidata [3] model and tool occlusion binary masks using a pre-trained SAM2 [13] model. We measure the surgical scene reconstruction capabilities of SGS using three standard metrics: LPIPS, PSNR, and SSIM. Our method achieves state-of-the-art (SOTA) results on all metrics.
Implementation details. All experiments are carried out on an NVIDIA RTX A6000, using the PyTorch framework. Adam [9] optimizer is used during training. The loss weights in Eq. 6 were set empirically to: $\lambda_{\text{1}}=0.7,\ \lambda_{\text{2}}=0.006,\ \lambda_{\text{3}}=1.0,\ \lambda_{\text{4}}=200.0,\ \lambda_{\text{5}}=1.0,\ \lambda_{\text{6}}=0.0001.$
EndoNeRF EndoGaussian LerPlane SG Ours Reference
Figure 3: Qualitative comparisons of our method (SGS) with EndoNeRF [16], EndoSurf [20], LerPlane [18], EndoGaussian [10], and SG [17].
4.1 Main Results and Ablation
We compare SGS against existing methods for real-time rendering of surgical scenes in the EndoNeRF [16] and StereoMIS [5] datasets. SDF-based approaches such as EndoSurf [20] require separate MLPs for deformation, geometry, and appearance, which increases computational complexity and limits real-time performance. Neural LerPlane [18] factorizes 4D scenes into 2D planes to reduce overhead, but struggles to capture complex deformations. EndoGaussian [10] shifts towards explicit depth-aware point clouds but relies heavily on pre-initialized geometry, making it less adaptable to dynamic changes. Lastly, SurgicalGaussian [17] introduces a deformation network that decouples motion and geometry, but struggles to render fine details in the surgical scene. SGS outperforms these previous methods in all datasets, achieving the lowest LPIPS, the highest PSNR and the highest SSIM, while qualitatively capturing specularity and anatomical features more accurately, as shown in Tab. 1 and Fig. 3.
@@ 4.1 Main Results and Ablation
Ablation study. We perform ablation studies in a "leave-one-out" fashion on the EndoNeRF [20] dataset. To assess the impact of our PIMI Gaussian initialization, we replace it with the prior SOTA GIDM [17] point cloud initialization. We also provide test-time results without perceptual, photometric, and TV loss. The results are summarized in Tab. 2. Homodirectional view-space positional gradients [19] are used during Gaussian Surfel densification for all experiments.
@@ 5 Conclusion
In this article, we introduce Surgical Gaussian Surfels (SGS), an explicit rendering method that takes advantage of the accurate geometry reconstruction capabilities of surface-aligned anisotropic splats. We utilize scale-constrained elliptical primitives to produce high-fidelity renders of surgical scenes. We define surface normals as the direction of the steepest density change, thereby generating accurate surgical tissue surface normals without normal priors. By leveraging both Projection-Based Iterative Mask Integration (PIMI) for robust point-cloud initialization and Gaussian Surfel densification based on homodirectional view-space positional gradients, our method achieves state-of-the-art reconstruction quality with real-time rendering performance. SGS has the potential to advance surgical scene understanding and support a range of intraoperative applications, including surgical navigation and robot-aided surgical procedures.

### v2

@@ 3.3 PIMI Gaussian Surfel Primitive Initialization
To mitigate these issues, we propose an initialization method, termed Projection-Based Iterative Mask Integration (PIMI) (Fig. 2(a)), which improves the quality and completeness of the initial point cloud by leveraging confidence-weighted depth aggregation and aggregated color mapping. The depth map $D_{t}$ for each frame is pre-processed to create a confidence-weighted depth map $D^{*}$ using a thresholded confidence mask, which ensures the exclusion of noisy depth data. Simultaneously, we aggregate the color maps $C_{t}$ to compute the refined color map $C^{*}$. This aggregation scheme at timestep $t$ is defined as:
@@
[…]
@@
Figure 4: Quantitative comparison of FFD-MLP, CutlassMLP [21], and Vanilla MLP on the EndoNerf [27] pulling dataset. Our implementation for deformable tissue prediction is, at times, 5× faster than other MLP implementations at inference time.
3.4 Fully Fused Deformation MLP (FFD-MLP)
For our tissue deformation network in Fig. 2(b), we aim to develop an architecture that is both fast and accurate in predicting motion fields. Inspired by [26] and [19], we devise a fully fused implementation of the MLP on modern GPUs to predict surgical scene deformation. We refer to our network as the Fully Fused Deformation MLP (FFD-MLP), a high-performance implementation of the multilayer perceptron designed explicitly for fast deformation network training within the SGS pipeline.
Built on Tiny-CUDA-NN [18], FFD-MLP achieves a significant speedup through GPU-optimized kernels that fuse multiple operations into a single CUDA kernel, eliminating memory bandwidth bottlenecks and maximizing computational throughput. The network architecture consists of 11 fully connected hidden layers, each with 128 neurons, processing 72-dimensional encoded features (63 spatial and 9 temporal) to produce a 10-dimensional deformation output (3, 4, and 3 for the position, rotation, and scaling of each Gaussian surfel, respectively).
The network employs a composite frequency encoding that matches the positional encoding used in standard NeRF-style [17] networks: $\gamma(\mathbf{p},t)=[\gamma_{xyz}(\mathbf{p}),\gamma_{t}(t)]$, where $\mathbf{p}\in\mathbb{R}^{3}$ represents 3D coordinates and $t\in\mathbb{R}$ represents time. The spatial encoding $\gamma_{xyz}$ uses 10 frequency bands: $\gamma_{xyz}(\mathbf{p})=[\mathbf{p},\sin(2^{0}\pi\mathbf{p}),\cos(2^{0}\pi\mathbf{p}),\ldots,\sin(2^{9}\pi\mathbf{p}),\cos(2^{9}\pi\mathbf{p})]$, while the temporal encoding $\gamma_{t}$ uses 4 frequency bands: $\gamma_{t}(t)=[t,\sin(2^{0}\pi t),\cos(2^{0}\pi t),\ldots,\sin(2^{3}\pi t),\cos(2^{3}\pi t)]$. This encoding provides multi-scale frequency information, enabling the network to learn both fine-grained local deformations and coarse global transformations. The core innovation lies in the fully fused CUDA implementation, which leverages NVIDIA’s Tensor Cores through Warp Matrix Multiply Accumulate (WMMA) [20] operations, processing 16×16×16 matrix multiplications in a single instruction while storing activations in shared memory and using register-based storage for weight matrices.
Empirical results demonstrate significant performance improvements—2.5$\times$ faster training, 5$\times$ higher iteration throughput (Fig. 4), and a 40% reduction in GPU memory—while maintaining superior rendering quality. Given
$(\delta\mathbf{x},\delta\mathbf{s},\delta\mathbf{q})=\text{FFD-MLP}\!\left(\gamma(\mathbf{p},t)\right),$ (11)
The predicted translation $\delta\mathbf{x}$, scale $\delta\mathbf{s}$, and rotation $\delta\mathbf{q}$ update the canonical parameters of each surfel:
$\mathbf{p}^{\prime}=\mathbf{p}+\delta\mathbf{x},\quad\mathbf{s}^{\prime}=\mathbf{s}\odot\delta\mathbf{s},\quad\mathbf{q}^{\prime}=\mathbf{q}\otimes\delta\mathbf{q}.$ (12)
where $\odot$ denotes element‑wise scaling and $\otimes$ quaternion composition.
@@ 3.5 Optimization
To capture fine surgical tissue detail and preserve geometric surfel–surface alignment, we optimize a weighted objective (see Fig. 2(c)) combining photometric fidelity, structural similarity, spatial smoothing, deformation regularization, perceptual alignment (DINO [2]), depth supervision, and normal consistency. The total objective is
$L_{\text{total}}=L_{\text{photo}}+\lambda_{\text{smooth}}L_{\text{TV}}+\lambda_{\text{pos}}L_{\text{pos}}+\lambda_{\text{cov}}L_{\text{cov}}+\\[-2.0pt] \lambda_{\text{per}}L_{\text{per}}+\lambda_{\text{normal}}L_{\text{n}}+\lambda_{\text{depth}}L_{\text{depth}}.$ (13)
Image Reconstruction (Photometric + SSIM + TV).
Pixel‑wise fidelity is enforced with an $L_{1}$ + SSIM blend. We penalize discrepancies in pixel intensity while accounting for surgical tool-masked regions:
$L_{\text{photo}}=(1-\lambda_{\text{SSIM}})\|I-\hat{I}\|_{1}\\[-2.0pt] {}+\lambda_{\text{SSIM}}\bigl{(}1-\text{SSIM}(I,\hat{I})\bigr{)},$ (14)
where $I$ and $\hat{I}$ are ground‑truth and rendered images. Spatial pixel smoothness is encouraged with total variation:
$L_{\text{TV}}=\sum_{p,q}\Bigl{(}\|\hat{I}_{p,q}-\hat{I}_{p-1,q}\|_{1}+\|\hat{I}_{p,q}-\hat{I}_{p,q-1}\|_{1}\Bigr{)}.$ (15)
Deformation Regularization (Position + Scale).
We encourage locally coherent motion and scale evolution across neighboring surfels $\mathcal{N}(i)$:
$L_{\text{pos}}=\sum_{i}\sum_{j\in\mathcal{N}(i)}\bigl{\|}x_{i}-x_{j}\bigr{\|}_{1}-\bigl{\|}\hat{x}_{i}-\hat{x}_{j}\bigr{\|}_{1},$ (16)
$L_{\text{cov}}=\sum_{i}\sum_{j\in\mathcal{N}(i)}\bigl{\|}s_{i}-s_{j}\bigr{\|}_{1}-\bigl{\|}\hat{s}_{i}-\hat{s}_{j}\bigr{\|}_{1},$ (17)
where $x_{i}$ and $x_{j}$ are the positions of neighboring surfels in the initial state, and $\hat{x}_{i}$ and $\hat{x}_{j}$ are their positions after deformation. In addition, $s_{i}$ and $s_{j}$ represent the scaling factors of neighboring surfels.
Perceptual (DINO) Loss.
High‑level appearance is matched using ViT‑based DINO features:
$L_{\text{per}}=1-\cos\!\bigl{(}\phi(\hat{I}),\phi(I)\bigr{)},$ (18)
where $\phi$ represents the DINO [2] multiscale semantic embedding extractor.
Depth Supervision.
Rendered depth is nudged toward sharp, mask‑filtered supervision using:
$L_{\text{depth}}=\|D_{\text{prediction}}-D_{\text{GT}}\|_{1}.$ (19)
Normal Consistency.
We encourage agreement between rendered normals and expected geometric normals from depth signals via a two‑target cosine ranking:
$L_{\text{n}}=\alpha\bigl{(}1-\cos(\mathbf{n}_{\text{rendered}},\mathbf{n}_{\text{median}})\bigr{)}\\[-2.0pt] {}+\beta\bigl{(}1-\cos(\mathbf{n}_{\text{rendered}},\mathbf{n}_{\text{expected}})\bigr{)},$ (20)
where $\mathbf{n}_{\text{median}}$ and $\mathbf{n}_{\text{expected}}$ are derived from scene geometry along the viewing ray; $\alpha=0.6$, $\beta=0.4$ in all experiments.
4 Experiments
Table 1: Quantitative comparisons of our method SGS with EndoNeRF [27], EndoSurf [33], LerPlane [30], EndoGaussian [15], SurgicalGaussian (SG) [29], EndoGS [34], and Deform3DGS [31].
Dataset Methods ”pulling” ”cutting” Rendering
@@ 4 Experiments
EndoGS 0.053 37.240 0.961 0.043 35.696 0.955 55 3 GB
Deform3DGS 0.063 38.367 0.961 0.042 38.771 0.956 225 5 GB
Ours 0.027 39.059 0.971 0.027 37.609 0.965 310 2 GB
”intestine” ”liver” FPS↑ GPU↓
@@
EndoGS 0.060 35.239 0.934 0.051 34.540 0.932 61 2 GB
Deform3DGS 0.131 32.589 0.897 0.110 32.929 0.900 196 5 GB
Ours 0.048 34.669 0.936 0.047 34.390 0.934 423 2 GB
4.1 Experiment settings
4.1.1 Datasets and evaluation.
We evaluate our method on two publicly available datasets, EndoNerf [27] and StereoMIS [9]. The EndoNerf dataset captures two cases of an in-vivo da Vinci robot prostatectomy from a single panoptic viewpoint. Each video comprises complex surgical tissue deformation and surgical tool occlusion. We also evaluate on two $\sim$5s video segments from the videos P2_7 and P2_8 of the StereoMIS dataset, which account for the breathing of porcine subjects and tissue deformation during a liver and intestinal procedure. Both datasets are sampled at a frame rate of 30 fps. We follow [33] and divide the training and test sets in a 7:1 ratio. We estimate scene depth on the StereoMIS dataset using a pre-trained Omnidata [7] model and surgical tool occlusion binary masks generated by a pre-trained SAM 2 [23] model. We measure the surgical scene reconstruction capabilities of SGS using three standard metrics: LPIPS, PSNR, and SSIM. We achieve state-of-the-art (SOTA) results on all metrics. We clearly show our qualitative and quantitative results against prior works in Fig. LABEL:fig:qualitative_comparison and Table 1, respectively.
4.1.2 Implementation Details.
All experiments are carried out on an NVIDIA RTX A6000 using the PyTorch framework. Adam [13] optimizer is used during training. The loss weights in Eq. 13 are set to:
$\lambda_{SSIM}=0.2,\quad\lambda_{smooth}=0.006,\quad\lambda_{pos}=1.0,\\ \quad\lambda_{cov}=200.0,\quad\lambda_{per}=1.0,\quad\lambda_{depth}=0.0001.$ (21)
4.2 Main Results
We compare SGS with existing methods for real-time rendering of surgical scenes in the EndoNeRF [27] and StereoMIS [9] datasets. SDF-based approaches such as EndoSurf [33] require separate MLPs for deformation, geometry, and appearance, which increases computational complexity and limits real-time performance. Neural LerPlane [30] factorizes 4D scenes into 2D planes to reduce overhead, but struggles to capture complex deformations. EndoGaussian [15] shifts towards explicit depth-aware point clouds but relies heavily on pre-initialized geometry, making it less adaptable to dynamic changes.
Furthermore, EndoGS [10] represents a surgical scene as a 4D volume, leveraging a multi-resolution HexPlane [1] to encode both spatial and temporal dynamics. Deform3DGS [31] enhances tissue deformation modeling by incorporating learnable mean and variance operators in Fourier and polynomial basis functions. Both methods face challenges relating to the complexity of representing highly dynamic and non-rigid tissue deformations, as well as the computational cost of handling multi-resolution representations and learnable basis functions, which may limit real-time applicability and efficiency in practical surgical scenarios.
Lastly, SurgicalGaussian [29] introduces a deformation network that decouples motion and geometry, but struggles to render fine details in the surgical scene. SGS mostly outperforms these previous methods in all datasets, achieving the lowest LPIPS, the highest PSNR, and the highest SSIM, while qualitatively capturing specularity and anatomical features more accurately, as shown in Fig. LABEL:fig:qualitative_comparison and Table 1, respectively.
4.2.1 Ablation Study
@@
We perform ablation studies in a ”leave-one-out” fashion on the EndoNerf [33] dataset. To assess the impact of our PIMI Gaussian initialization scheme, we replace it with the GIDM [29] point cloud initialization. We also provide test-time results without perceptual, photometric, and TV loss. The results are summarized in Table 2. The similarity in the ablation results to the full model could be attributed to the homodirectional view-space positional gradient guidance [32] during Gaussian surfel densification at training time. We also conduct an ablation study on the surgical scene deformation prediction network (Fig. 2(b)) within our rendering pipeline. We replace our FFD-MLP (sec. 3.4) with the CutlassMLP [21] during both training and inference times, along with a separate training and inference run with a ”vanilla” MLP implementation. As shown in Fig. 4, our FFD-MLP attains inference speeds 5× faster than both the CutlassMLP and classical MLP implementations.
4.2.2 Comparison on EndoNerf Dataset
The EndoNerf dataset [27] is a mainstay in the surgical scene reconstruction domain. The dataset comprises six curated clips from DaVinci robotic prostatectomy procedures. Each clip spans 4$\sim$8 seconds and is captured from a fixed stereo camera viewpoint. The dataset features complex surgical scenes with non-rigid tissue deformation and frequent tool occlusion. We qualitatively and quantitatively compare our method using the EndoNerf dataset against SOTA methods [27, 33, 30, 15, 29, 34, 31] and highlight our results in Fig. LABEL:fig:qualitative_comparison and Table 1, respectively. We train and test our method on tissue pulling and tissue cutting videos, where it achieves lower perceptual error (LPIPS), higher signal fidelity (PSNR), and stronger structural consistency (SSIM). This indicates that SGS renders dynamic tissue motion fields closer to the ground-truth surgical tissue motion and photometric features at the pixel level. As a result of our efficiency gains from FFD-MLP (sec. 3.4), this improvement in rendering quality does not come at the expense of rendering efficiency and minimal GPU memory usage, demonstrating the practical deployability of SGS in resource-constrained surgical robotics settings.
4.2.3 Comparison on StereoMIS Dataset
The StereoMIS dataset [9] comprises 11 in‑vivo porcine sequences captured at 15fps by the DaVinci Xi stereo endoscope. The dataset captures complex physiological motion of surgical tissue undergoing breathing-induced deformation, instrument occlusion, and manipulation. We benchmark SGS against state-of-the-art methods [27, 33, 30, 15, 29, 34, 31] in representative intestine and liver sequences (Table 1), achieving the lowest perceptual error (LPIPS) and the highest structural similarity (SSIM) in both scenes, demonstrating that SGS more accurately reconstructs fine-grained tissue appearance and geometry. EndoGS [34], which models the scene as a time-varying 4D volume via multi-resolution HexPlane [1] encoding, yields marginally higher PSNR in the liver case, indicating its strength in signal fidelity when leveraging high-capacity spatiotemporal representations. Crucially, SGS matches or exceeds the overall reconstruction quality of EndoGS while delivering higher frame rates and maintaining a smaller GPU footprint.
@@ 5 Conclusion
In this article, we introduce Surgical Gaussian Surfels (SGS), an explicit rendering method that takes advantage of the accurate geometry reconstruction capabilities of surface-aligned anisotropic splats. We utilize scale-constrained elliptical primitives to produce high-fidelity renders of surgical scenes. We define surface normals as the direction of the steepest density change, thereby generating accurate surgical tissue surface normals without normal priors. In addition, we introduce FFD-MLP, an implementation of the tiny-cuda-nn for extremely fast surgical tissue deformation field prediction. Furthermore, by leveraging both Projection-Based Iterative Mask Integration (PIMI) for robust point-cloud initialization and Gaussian surfel densification based on homodirectional view-space positional gradients, our method achieves state-of-the-art reconstruction quality with real-time rendering performance. A limitation of our current implementation is that training, although on the order of minutes, cannot be done on the fly; a common setback of volumetric rendering and point primitive methods. SGS has the potential to advance surgical scene understanding and support a range of intraoperative applications, including surgical navigation and robot-aided surgical procedures.